OpenStudy (koikkara):

If limit(x=0), then sin x/x=?

OpenStudy (dls):

1

hartnn (hartnn):

if \(x\rightarrow 0 ,then , \frac{\sin x}{x}=1 \) because \(\lim \limits_{x \rightarrow0}\frac{\sin x}{x}=1\)

OpenStudy (dls):

haha student outrules his teacher xD

hartnn (hartnn):

i am not in any kind of race :P

OpenStudy (dls):

hmm..@Koikkara got it?

OpenStudy (koikkara):

hey guys, thanks n specially appreciate u, two...for the tip...!!

OpenStudy (shubhamsrg):

well.. sin x/x = sin1 //isnt it? ;)

OpenStudy (koikkara):

Something new......... Just use hopital's rule for that ... sin(x) / x == cos(x) /1 and cos(x) for x->0 = 1

Parth (parthkohli):

@Koikkara: It'd be fun to know how you prove \(\dfrac{d}{dx}\sin(x) = \cos(x)\)

OpenStudy (shubhamsrg):

taylor series hmm ?

hartnn (hartnn):

limit definitaion would do it....

OpenStudy (shubhamsrg):

infact, taylor series itself makes use of the face that sin'x = cosx :P

Parth (parthkohli):

Also\[\dfrac{\sin(x)}{x} \ne \dfrac{\cos(x)}{1}\]@Koikkara :P

Parth (parthkohli):

@shubhamsrg Haha yes, the expansions would work =)

OpenStudy (koikkara):

According to L'Hopital's rule, it is circular reasoning for this limit. L'Hopital's rule depends on the knowing that the derivative of sin(x) is cos(x), and in order to have that, we need to know that lim_{x->0} sin(x)/x = 1.

OpenStudy (hba):

You can use FPM.

Parth (parthkohli):

Prove the L'Hopital's Rule.

OpenStudy (shubhamsrg):

i wont say use of expansions would be good.. since expansion itself uses the fact that sinx ' = cos x

OpenStudy (shubhamsrg):

as @hartnn said,,definiton of limits will help our cause..

Parth (parthkohli):

Maybe term-by-term derivative? @shubhamsrg

hartnn (hartnn):

use limit definition to prove that derivative....

Parth (parthkohli):

Oh yes.

Parth (parthkohli):

\[\frac{\sin(x +h) -\sin(x)}{h}\]Hmm. \(\sin(a + b)\) formula and stuff

hartnn (hartnn):

or sin C-sin D

OpenStudy (shubhamsrg):

leme try geometrically //hmmm  sin x = a/c we need to find rate of change of sinx if x changes hmm..